# spacing=nonfrench

### v2

@@ Abstract
We present a theoretical study of 1D chemical systems subjected to temperature gradients. A complete theoretical framework describes the behavior of the system induced by various temperature profiles. An exact mathematical derivation was established for a simple two-compartment model and was generalized to arbitrary reaction-diffusion systems based on numerical models. An experimental system was eventually scaled and tuned to optimize either nonequilibrium chemical transport or reaction.
@@ I Introduction
The dynamic behavior of the nonequilibrium steady states is theoretically understood[6, 7, 8]. In contrast, the issue of steadily sustaining an nonequilibrium reactive state has received much less attention and remains a major challenge[9]. In practice, its realization requires either introducing constraints to prevent the effective relaxation of the reactive processes towards equilibrium[10, 11, 12], or steadily applying a periodic excitation on the reactive system.[13] We previously demonstrated the relevance of the first approach by frustrating the relaxation of reactive processes by the interference of molecular motion[14, 15], a strategy to generate reaction-diffusion cycles that are thought to establish spatial gradients of signaling activities in living cells[16, 17]. However, the efficiency of this first approach has not been theoretically analyzed, and is the object of the present manuscript. xpired. In our previous work[14, 15], we used light as an energy source to maintain a reactive system in a nonequilibrium state. In the present study, we favored a more general approach and adopted the application of a temperature gradient. Despite its chemical relevance, the behavior of reactive media in contact with two thermostats at different temperatures has included glorious names (including Nernst[18], Dirac[19], Prigogine[20, 21]). Several papers have been published since then, but have been mainly concerned with the gas phase and the question of how a chemical reaction that occurs in a system modifies its thermal conductivity[22, 23, 24, 25, 26, 27, 28, 29, 30]. Recently, heat flow has been applied to generate proton gradients and pH oscillations in microscale aqueous solutions[31, 32]. In the latter works, the reactive system was often assumed to be at chemical equilibrium within the temperature gradient.
In this paper, we level off this assumption, which has been predicted to generate attractive behavior such as dissipation-driven selection of states in nonequilibrium chemical networks[33], or emergent thermophoretic behavior in chemical reaction systems[34]. More specifically, we identify the optimal regime for shifting away a reactive system from chemical equilibrium at a steady state. We analyze optimization of the flux of the reaction and diffusion processes and of the corresponding entropy production[35], which are key elements in the propagation of energy into a set of coupled reactions. To achieve this optimization, we benefit from a fruitful analogy between entropy production by a chemical reaction and the ohm dissipated power by electrical resistors. We then identify the conditions for the spatial localization or delocalization of nonequilibrium activity. Finally, we assessed the possibility of propagating this primary energy transfer from the heat sources and sinks to secondary athermic reaction or transport processes.
@@ II.1 Chemical model
The studied chemical system was reduced to a single isomerization reaction
@@ II.2 Chemical response to a temperature gradient between two homogeneous compartments
The general behavior of a reaction driven by temperature gradients was first characterized in the limits of two compartments, each of which was homogeneous in temperature and concentration. This corresponds to two stirred systems; each compartment $A/B$ is connected to the other by chemical exchange of the reaction rate constants $\delta_{i}$ for each compound of concentration $u_{i}$.
This setup can be reduced to a set of ordinary differential equations with a set of parameters for each compartment. Analytical solutions for the resulting steady state were obtained. A second-order Taylor development as a function of the temperature gradient $\theta_{g}$ leads to simpler but more informative solutions.
II.2.1 Parameters
The general properties of the system are described based on its characteristic dimensions: concentration $c_{0}$ (corresponding to the total concentration), temperature $T_{0}$ (corresponding to the median system temperature) and time $t_{0}$. All subsequent system parameters were nondimensionalized using these characteristic parameters (see Table 1).
@@ II.2.1 Parameters
The system is thus described in the nondimensionalized concentrations $u_{i,j}$ of compound $i$ in compartment $j$, relative temperature deviation $\theta$, and time $\bar{t}$. Each compartment is thermostated at temperature $\theta_{a}=-\theta_{g}$ and $\theta_{b}=+\theta_{g}$, with $\theta_{g}\in[0,1]$; the limit cases correspond, respectively, to a uniform temperature $T_{0}$ and a temperature gradient from 00 K to $2T_{0}$; the median temperature is $T_{0}$ in all situations.
@@
  • $\kappa$ is the square root of the thermodynamic constant at $T_{0}$, which reports the reaction asymmetry,
@@
  • $\gamma$, which reports the asymmetry of the activation energy profile.
@@
  • $\lambda$ is the squared root of the exchange rate constant, which reports the asymmetry between the exchange rate of the two compounds (see Eqs. (25af)).
All kinetic and thermodynamic constants can be expressed as a function of these parameters at any point in the system, based on the Arrhenius equation (see Eqs. (25aea)-(25aec)). $\kappa$ and $\lambda$ can be combined into two secondary parameters $\kappa^{\prime}$ and $\lambda^{\prime}$, leading to simpler mathematical representations. They represent the reaction and exchange asymmetry, respectively, with a maximal value of $1$ for a perfectly symmetric process and 00 for a process totally displaced in a given direction.
The resulting system behavior in the steady state was then characterized in terms of three dimensionless parameters: chemical force $\alpha$, chemical flux $\varphi$, and entropic dissipation $\sigma$, which can be expressed as:
$\sigma$ $=\alpha\varphi$ (2)
II.2.2 Establishment of a nonequilibrium steady state
@@
This results in a nonequilibrium state, whose distance from equilibrium can be quantified by the chemical force $\alpha$, as the sum of the chemical forces of each chemical reaction in each compartment and the chemical forces of each chemical exchange between the two compartments. It can then be expressed for low values of $\theta_{g}$ (see Eqs. (25aiar)-(25aias)) as:
$\alpha$ $=2\eta\theta_{g}$ (3)
@@
In response to this non-zero chemical force, a circular reaction-diffusion flux is established. Chemical reaction fluxes are processed in opposite directions in each compartment, compensated by the continuous exchange of each compound in the opposite direction (see Fig. 2A). By analogy to electric circuits, $\alpha$ can be interpreted as the potential difference and $\varphi$ as the intensity. These are expressed as (see Eqs. (25aiba) and (25aibd)):
$\varphi$ $=\frac{\alpha}{R_{\text{tot}}}$ (4)
@@
$R_{\text{tot}}$ $=\frac{4}{\kappa^{\prime}}\left(\frac{1}{\rho}+\frac{1}{\lambda^{\prime}\delta}\right)$ (5)
This chemical resistance $R_{\text{tot}}$ is decomposed as the sum of two terms: the first one is directly linked to the chemical reaction (term in $\nicefrac{{1}}{{\rho}}$), and the second one to the chemical exchanges (term in $\nicefrac{{1}}{{\lambda^{\prime}\delta}}$).
Figure 2: Establishment of a nonequilibrium steady reaction-diffusion cycle driven by a temperature gradient. Cyclic flux $\varphi$ in response to nonzero chemical force $\alpha$ (A); distribution of the entropy dissipation in each chemical reaction or exchange process (B).
II.2.3 Process Distribution
This steady nonequilibrium reaction-diffusion flux leads to a continuous dissipative process characterized by $\sigma$, as an analog of electric power dissipation[36] according to Eq. 2. The energy extracted from the temperature gradient is dissipated differently in each process (see Fig. 2B), and can be decomposed as a sum of terms specific to each process, according to each chemical resistance as (see Eqs. (25aian)-(25aiap)):
$\sigma$ $=\sum_{i}\sigma_{i}=\sum_{i}R_{i}\varphi^{2}$ (6)
$R_{\text{tot}}$ $=\sum_{i}R_{i}$ (7)
Reactions and exchanges
The full dissipation $\sigma$ is first dispatched between the reaction term $\sigma_{r}$ and the exchange term $\sigma_{d}$:
$\sigma_{r}=\frac{\lambda^{\prime}\delta}{\lambda^{\prime}\delta+\rho}\sigma\qquad$ $;\quad\sigma_{d}=\frac{\rho}{\lambda^{\prime}\delta+\rho}\sigma$ (8)
with
$R_{r}=\frac{4}{\kappa^{\prime}\rho}\qquad$ $;\quad R_{d}=\frac{4}{\kappa^{\prime}\lambda^{\prime}\delta}$ (9)
@@ II.2.3 Process Distribution
Reaction balance between compartments
The reaction dissipation is decomposed in each compartment as:
$\sigma_{ra}$ $=\frac{1+\varepsilon\theta_{g}}{2}\sigma_{r}$ (10)
$\sigma_{rb}$ $=\frac{1-\varepsilon\theta_{g}}{2}\sigma_{r}$ (11)
The parameter $\varepsilon$ (see Eqs. (25aiaw)) comes down to an apparent activation energy for the full system, varying from $\varepsilon_{+}$ to $\varepsilon_{-}$ through the average $\beta$ value, depending on the global symmetry of the system: it tends to $\beta$ when $\kappa\lambda\sim 1$, to $\beta(1-\gamma)=\varepsilon_{+}$ when $\kappa\lambda\gg 1$, and to $\beta(1+\gamma)=\varepsilon_{-}$ when $\kappa\lambda\ll 1$.
@@
The exchange dissipation can be decomposed by the exchange of each compound as:
$\sigma_{d1}$ $=\frac{\lambda^{2}\kappa^{2}}{1+\lambda^{2}\kappa^{2}}\sigma_{d}$ (12)
$\sigma_{d2}$ $=\frac{1}{1+\lambda^{2}\kappa^{2}}\sigma_{d}$ (13)
It is symmetrically distributed between the two compounds when $\lambda\kappa=1$. The asymmetrization of this exchange is quantified by $\lambda\kappa$, with a $\sigma_{d}\approx\sigma_{d1}$ for $\lambda\kappa\gg 1$ (the exchange dissipation is then governed by $U_{1}$), and with a $\sigma_{d}\approx\sigma_{d2}$ for $\lambda\kappa\ll 1$ (the exchange dissipation is then governed by $U_{2}$).
II.2.4 Optimal conditions
To maximize the efficiency of the nonequilibrium process, it is first necessary to maximize its distance from the equilibrium, i.e., to maximize $\alpha$. According to Eq. (3), this implies maximization of both $\theta_{g}$ and $\eta$. Thus, high temperature gradients coupled with a high-enthalpy chemical reaction should be favored as a necessary condition.
Moreover, it is necessary to consider the dynamics of the system to avoid a kinetically locked nonequilibrium state. Thus, one should also maximize the resulting flux $\varphi$, i.e. the intensity at which the chemical reaction is processed, as the quantity of matter that is converted at each moment in each compartment, and the quantity of matter that is transferred at each moment from one compartment to another.
From Eqs. (4)-(5), maximizing the steady nonequilibrium flux $\varphi=\nicefrac{{\alpha}}{{R_{\text{tot}}}}$ implies the minimization of its chemical resistance $R_{\text{tot}}$, in addition to $\alpha$ maximization. This implies:
  • to maximize parameter $\kappa^{\prime}$; its largest possible value is $\kappa^{\prime}=1$, corresponding to a chemical reaction perfectly balanced between its reactants at $T_{0}$ (i.e. $K_{0}=1$),
  • to maximize parameter $\lambda^{\prime}$ to $1$, corresponding to symmetric matter exchange (see Eqs. (12)-(13)); this implies that an imbalance in the chemical equilibrium (characterized by $\kappa$) can be compensated for by an imbalance in the chemical exchange (characterized by $\lambda$),

### v3

@@ Abstract
We present a theoretical study of 1D chemical systems subjected to temperature gradients, for sustaining nonequilibrium chemical fluxes. A complete theoretical framework describes the behavior of the system induced by various temperature profiles. An exact mathematical derivation was established for a simple two-compartment model and was generalized to arbitrary reaction-diffusion systems based on numerical models. An experimental system was eventually scaled and tuned to optimize either nonequilibrium chemical transport or reaction.
@@ I Introduction
The logic of preparative chemistry is in sharp contrast with the one of biochemical systems. The former focuses primarily on multistep synthesis, with the objective of producing pure compounds in optimal yield through individually optimized reactions. In living cells, the reactions are also grouped in sets. However, they occur in cycles, like the Krebs and Calvin cycles, where reactants and products are not distinguished from each other, but simultaneously coexist in non-equilibrium steady states. Hence, although sharing a same common canon of reactivity, a cell operates differently than a chemical reactor seeking maximal yields. What advantages does this mode of operation afford?
@@
The dynamic behavior of the nonequilibrium steady states is theoretically understood [6, 7, 8]. In contrast, the issue of steadily sustaining an nonequilibrium reactive state has received much less attention and remains a major challenge [9]. In practice, its realization requires either introducing constraints to prevent the effective relaxation of the reactive processes towards equilibrium [10, 11, 12], or steadily applying a periodic excitation on the reactive system. [13] We previously demonstrated the relevance of the first approach by frustrating the relaxation of reactive processes by the interference of molecular motion [14, 15], a strategy to generate reaction-diffusion cycles that are thought to establish spatial gradients of signaling activities in living cells [16, 17]. However, the efficiency of this first approach has not been theoretically analyzed, and is the object of the present manuscript.
In our previous work [14, 15], we used light as an energy source to maintain a reactive system in a nonequilibrium state. In the present study, we favored a more general approach and adopted the application of a temperature gradient. Despite its chemical relevance, the behavior of reactive media in contact with two thermostats at different temperatures has included glorious names (including Nernst [18], Dirac [19], Prigogine [20, 21]). Several papers have been published since then, but have been mainly concerned with the gas phase and the question of how a chemical reaction that occurs in a system modifies its thermal conductivity [22, 23, 24, 25, 26, 27, 28, 29, 30]. Recently, heat flow has been applied to generate proton gradients and pH oscillations in microscale aqueous solutions [31, 32]. In the latter works, the reactive system was often assumed to be at chemical equilibrium within the temperature gradient.
In this paper, we level off this assumption, which has been predicted to generate attractive behavior such as dissipation-driven selection of states in nonequilibrium chemical networks [33], or emergent thermophoretic behavior in chemical reaction systems [34]. We investigate the behavior of an active chemical system submitted to a nonequilibrium force, induced by an externally imposed temperature gradient. The primary objective is to determine the optimal characteristics and conditions that will lead to maximizing an internal nonequilibrium reaction-diffusion flux within the system. This consists in a simple chemical reaction subjected to an inhomogeneous field of temperature. The associated temperature gradient directly generates a large entropy production by heat diffusion; the aim is to redirect a maximal part of this source of nonequilibrium to chemical processes.
In this context, the temperature gradient-induced steady chemical force and the resulting steady chemical flux must both be maximized, with the ultimate objective of maximizing the entropy rate produced specifically by this internally sustained cyclic process. This objective was further refined by recognizing that the focus might entail optimizing either the transport process or the reaction process on their own, or balancing both processes. To achieve this optimization, we benefited from a fruitful analogy between entropy production by a chemical reaction and the ohm dissipated power by electrical resistors.
We then identified the conditions for the spatial localization or delocalization of nonequilibrium activity, emphasizing the important role of the system scaling and of the shape of the temperature gradient. Finally, we assessed the possibility of propagating this primary energy transfer from the heat sources and sinks to secondary athermic reaction or transport processes.
@@ II.1 Chemical model
The studied chemical system was reduced to experiencing a single isomerization reaction
@@
where the rate constants $k_{+}$ and $k_{-}$ are linked to the thermodynamic constant $K$ with
$K=\frac{k_{+}}{k_{-}}$ (2)
@@
Using thermodynamic and Arrhenius relationships, the variation in the constants with temperature can be expressed as:
$K$ $=K^{\infty}e^{-\frac{\Delta_{r}H_{0}}{RT}}$ (3a)
$k_{+}$ $=k_{+}^{\infty}e^{-\frac{E_{a,+}}{RT}}$ (3b)
$k_{-}$ $=k_{-}^{\infty}e^{-\frac{E_{a,-}}{RT}}$ (3c)
with
$\Delta_{r}H_{0}$ $=E_{a,+}-E_{a,-}$ (3d)
@@ II.2 Chemical response to a temperature gradient between two homogeneous compartments
The general behavior of a reaction driven by temperature gradients was first characterized in the limits of two compartments, each of which was homogeneous in temperature and concentration. This corresponds to two stirred systems; each compartment $A/B$ is connected to the other by chemical exchange with the reaction rate constant $k_{d,i}$ for each compound of concentration $c_{i}$:
@@
II.2.1 Parameters
The general properties of the system are described based on its characteristic dimensions: concentration $c_{0}$ (corresponding to the total concentration), temperature $T_{0}$ (corresponding to the median system temperature) and time $t_{0}$. All subsequent system parameters were nondimensionalized using these characteristic parameters (see Table 1). This leads to:
$K$ $=K_{0}e^{\frac{\Delta_{r}H_{0}}{RT_{0}}\left(\frac{T-T_{0}}{T}\right)}$ (4aca)
$k_{+}$ $=k_{+,0}e^{\frac{E_{a,+}}{RT_{0}}\left(\frac{T-T_{0}}{T}\right)}$ (4acb)
$k_{-}$ $=k_{-,0}e^{\frac{E_{a,-}}{RT_{0}}\left(\frac{T-T_{0}}{T}\right)}$ (4acc)
with $K_{0}=K(T_{0})$, $k_{+,0}=k_{+}(T_{0})$, $k_{-,0}=k_{-}(T_{0})$.
This is then nondimensionalized as:
$\bar{k}_{+}$ $=\kappa\rho e^{\beta(1-\gamma)\frac{\theta}{\theta+1}}$ (4ada)
$\bar{k}_{-}$ $=\kappa^{-1}\rho e^{\beta(1+\gamma)\frac{\theta}{\theta+1}}$ (4adb)
$K$ $=\kappa^{2}e^{2\beta\gamma\frac{\theta}{\theta+1}}$ (4adc)
$\eta$ $=2\beta\gamma$ (4add)
$\varepsilon_{+}$ $=\beta(1-\gamma)$ (4ade)
$\varepsilon_{-}$ $=\beta(1+\gamma)$ (4adf)
The diffusion is modeled as a first-order exchange reaction between two compartments as a function of the geometric mean of the exchange kinetic constant $\delta$:
$\delta_{1}$ $=\lambda\delta$ (4aea)
$\delta_{2}$ $=\lambda^{-1}\delta$ (4aeb)
The system can then be reduced to a set of ordinary differential equations with a set of nondimensionalized parameters:
$\frac{du_{1a}}{d\bar{t}}$ $=\lambda^{-1}\delta\left(-u_{1a}+u_{1b}\right)-\bar{k}_{+,a}u_{1a}+\bar{k}_{-,a}u_{2a}$ (4afa)
$=-\varphi_{d1}-\varphi_{ra}$ (4afb)
$\frac{du_{1b}}{d\bar{t}}$ $=\lambda^{-1}\delta\left(u_{1a}-u_{1b}\right)-\bar{k}_{+,b}u_{1b}+\bar{k}_{-,b}u_{2b}$ (4afc)
$=+\varphi_{d1}-\varphi_{rb}$ (4afd)
$\frac{du_{2a}}{d\bar{t}}$ $=\lambda\delta\left(-u_{2a}+u_{2b}\right)+\bar{k}_{+,a}u_{1a}-\bar{k}_{-,a}u_{2a}$ (4afe)
$=-\varphi_{d2}+\varphi_{ra}$ (4aff)
$\frac{du_{2b}}{d\bar{t}}$ $=\lambda\delta\left(u_{2a}-u_{2b}\right)+\bar{k}_{+,b}u_{1b}-\bar{k}_{-,b}u_{2b}$ (4afg)
$=+\varphi_{d2}+\varphi_{rb}$ (4afh)
The system is thus described in the nondimensionalized concentrations $u_{i,j}$ of compound $i$ in compartment $j$, relative temperature deviation $\theta$, and time $\bar{t}$. Each compartment is thermostated at temperature $\theta_{a}=-\theta_{g}$ and $\theta_{b}=+\theta_{g}$, with $\theta_{g}\in[0,1]$; the limit cases correspond, respectively, to a uniform temperature $T_{0}$ and a temperature gradient from 00 K to $2T_{0}$; the median temperature is $T_{0}$ in all situations. The initial concentration is $1$ in each compartment, i.e.:
$u_{1a}(0)+u_{2a}(0)$ $=u_{1b}(0)+u_{2b}(0)=1$ (4ag)
Analytical solutions for the resulting steady state were obtained. A second-order Taylor development as a function of the temperature gradient $\theta_{g}$ leads to simpler but more informative solutions (see Appendix C for details).
The resulting system behavior in the steady state was then characterized in terms of three dimensionless parameters: chemical force $\alpha$, chemical flux $\varphi$, and entropic production $\sigma$, which can be expressed as:
$\sigma$ $=\alpha\varphi$ (4ah)
II.2.2 Parameter description
@@
  • $\kappa$ is the square root of the thermodynamic constant at $T_{0}$, which reports the reaction asymmetry between the forward and backward reactions, with $\kappa=1$ for $u_{1}=u_{2}$ at equilibrium at $T_{0}$,
@@
  • $\gamma$, which reports the asymmetry of the activation energy profile (see Fig. S3 in SI for a graphical representation).
@@
  • $\lambda$ is the squared root of the exchange rate constant, which reports the asymmetry between the exchange rate of the two compounds (see Eqs. (4ae)).
  • $\kappa$ and $\lambda$ can be combined into two secondary parameters $\kappa^{\prime}$ and $\lambda^{\prime}$, leading to simpler mathematical representations. They represent the reaction and exchange asymmetry, respectively, with a maximal value of $1$ for a perfectly symmetric process and 00 for a process totally displaced in a given direction.
II.2.3 Establishment of a nonequilibrium steady state
@@
This results in a nonequilibrium state, whose distance from equilibrium can be quantified by the chemical force $\alpha$, as the sum of the chemical forces of each chemical reaction in each compartment and the chemical forces of each chemical exchange between the two compartments. It can then be expressed for low values of $\theta_{g}$ (see Eqs. (4bvbwci)-(4bvbwcj)) as:
$\alpha$ $=2\eta\theta_{g}$ (4ai)
@@
In response to this non-zero chemical force, a circular reaction-diffusion flux is established. Chemical reaction fluxes are processed in opposite directions in each compartment, compensated by the continuous exchange of each compound in the opposite direction (see Fig. 2A). By analogy to electric circuits, $\alpha$ can be interpreted as the potential difference and $\varphi$ as the intensity. These are expressed as (see Eqs. (4bvbwcr) and (4bvbwcu)):
$\varphi$ $=\frac{\alpha}{R_{\text{tot}}}$ (4aj)
In the case of the linear regime, characterized by low values of $\theta_{g}$, this chemical resistance can be expressed as:
$R_{\text{tot}}$ $=\frac{4}{\kappa^{\prime}}\left(\frac{1}{\rho}+\frac{1}{\lambda^{\prime}\delta}\right)$ (4ak)
It is decomposed as the sum of two terms: the first one is directly linked to the chemical reaction (term in $\nicefrac{{1}}{{\rho}}$), and the second one to the chemical exchanges (term in $\nicefrac{{1}}{{\lambda^{\prime}\delta}}$). Outside the linear regime, for large values of $\theta_{g}$, this chemical resistance will not be constant and may depend on the chemical force.
Figure 2: Establishment of a nonequilibrium steady reaction-diffusion cycle driven by a temperature gradient. Cyclic flux $\varphi$ in response to nonzero chemical force $\alpha$ (A); full entropy production $\sigma_{\mathrm{tot}}$ as a sum of thermal ($\sigma_{\mathrm{th}}$, see appendix B) and chemical ($\sigma$) contributions (B); distribution of the entropy production in each chemical reaction or exchange process (C).
II.2.4 Process Distribution
This steady nonequilibrium reaction-diffusion flux leads to a continuous dissipative process characterized by $\sigma$, as an analog of electric power dissipation [35] according to Eq. (4ah). The corresponding chemical entropy production taps on the important thermal entropy production by heat diffusion generated by the temperature gradient (Fig. 2B and appendix B). The energy extracted from the temperature gradient is dissipated differently in each process (see Fig. 2C), and can be decomposed as a sum of terms specific to each process, according to each chemical resistance as (see Eqs. (4bvbwce)-(4bvbwcg)):
$\sigma$ $=\sum_{i}\sigma_{i}=\sum_{i}R_{i}\varphi^{2}$ (4al)
$R_{\text{tot}}$ $=\sum_{i}R_{i}$ (4am)
Reactions and exchanges
The full entropy production $\sigma$ is first dispatched between the reaction term $\sigma_{r}$ and the exchange term $\sigma_{d}$:
$\sigma_{r}=\frac{\lambda^{\prime}\delta}{\lambda^{\prime}\delta+\rho}\sigma\qquad$ $;\quad\sigma_{d}=\frac{\rho}{\lambda^{\prime}\delta+\rho}\sigma$ (4an)
@@ II.2.4 Process Distribution
$R_{r}=\frac{4}{\kappa^{\prime}\rho}\qquad$ $;\quad R_{d}=\frac{4}{\kappa^{\prime}\lambda^{\prime}\delta}$ (4ao)
and
$\sigma=$ $\frac{\lambda^{\prime}\delta\rho}{\lambda^{\prime}\delta+\rho}\kappa^{\prime}\cdot\eta^{2}\theta_{g}^{2}$ (4ap)
@@
Reaction balance between the compartments
The reaction entropy production is decomposed in each compartment as:
$\sigma_{ra}$ $=\frac{1+\varepsilon\theta_{g}}{2}\sigma_{r}$ (4aq)
$\sigma_{rb}$ $=\frac{1-\varepsilon\theta_{g}}{2}\sigma_{r}$ (4ar)
The parameter $\varepsilon$ (see Eqs. (4bvbwcn)) comes down to an apparent activation energy for the full system, varying from $\varepsilon_{+}$ to $\varepsilon_{-}$ through the average $\beta$ value, depending on the global symmetry of the system: it tends to $\beta$ when $\kappa\lambda\sim 1$, to $\beta(1-\gamma)=\varepsilon_{+}$ when $\kappa\lambda\gg 1$, and to $\beta(1+\gamma)=\varepsilon_{-}$ when $\kappa\lambda\ll 1$.
@@
The exchange entropy production can be decomposed by the exchange of each compound as:
$\sigma_{d1}$ $=\frac{\lambda^{2}\kappa^{2}}{1+\lambda^{2}\kappa^{2}}\sigma_{d}$ (4as)
$\sigma_{d2}$ $=\frac{1}{1+\lambda^{2}\kappa^{2}}\sigma_{d}$ (4at)
It is symmetrically distributed between the two compounds when $\lambda\kappa=1$. The asymmetrization of this exchange is quantified by $\lambda\kappa$, with a $\sigma_{d}\approx\sigma_{d1}$ for $\lambda\kappa\gg 1$ (the exchange entropy production is then governed by $U_{1}$), and with a $\sigma_{d}\approx\sigma_{d2}$ for $\lambda\kappa\ll 1$ (the exchange entropy production is then governed by $U_{2}$).
II.2.5 Optimal conditions
Considering that the temperature gradient is fixed and cannot be optimized, several criteria can be used to evaluate the efficiency of the system to generate a steady nonequilibrium reaction-diffusion process. The chemical force $\alpha$ reflects how far the system is pushed away from the equilibrium state ($\alpha=0$ corresponding to the equilibrium state). The chemical flux $\varphi$ quantifies the system response as the intensity of the resulting reaction-diffusion flux from the chemical force, as the quantity of matter that is converted at each moment in each compartment, and the quantity of matter that is transferred at each moment from one compartment to another ($\varphi=0$ corresponding to a static, kinetically locked system). The rate of entropy production $\sigma$ combines both thermodynamic and kinetic aspects, and quantifies the nonequilibrium activity of the system.
At first, according to Eq. (4ai), the chemical force $\alpha$ is simply proportional to both $\theta_{g}$ and $\eta$; this entails that the temperature gradient shall be applied to a high-enthalpy chemical reaction as a first requirement from driving the system in a non-equilibrium state.
Secondly, from Eqs. (4aj)-(4ak), maximizing the steady nonequilibrium flux $\varphi=\nicefrac{{\alpha}}{{R_{\text{tot}}}}$ implies the minimization of its chemical resistance $R_{\text{tot}}$. This implies:
  • to maximize parameter $\kappa^{\prime}$; its largest possible value is $\kappa^{\prime}=1$, corresponding to a chemical reaction perfectly balanced between the backward and forward reactions at $T_{0}$ (i.e. $K_{0}=1$),
  • to maximize parameter $\lambda^{\prime}$ to $1$, corresponding to symmetric matter exchange (see Eqs. (4as)-(4at)); this implies that an imbalance in the chemical equilibrium at $T_{0}$ (characterized by $\kappa$) can be compensated for by an imbalance in the chemical exchange of each compound (characterized by $\lambda$),